a bag contains 20 red marbles, 16 green marbles, and 12 blue marbles. If you select one marble at random form the bah, what is the probability that the marble is green?
Answer by stanbon(75887)

a bag contains 20 red marbles, 16 green marbles, and 12 blue marbles. If you select one marble at random from the bag, what is the probability that the marble is green?
------------------------
# of ways to succeed: 16
# of possible outcomes: 48
-------------
P(one green) = 16/48 = 1/3
================================
Cheers,
Stan H.